OpenStudy (tylerd):

Write the equation of the ellipse 36 x^2 + 25 y^2 + 144x + 250y -131 = 0 in standard form

OpenStudy (tylerd):

\[36(x+2)^2+25(y+5)^2=131+144+625\]

OpenStudy (tylerd):

could this work? i wonder.

OpenStudy (tylerd):

36 times 25 = 900

OpenStudy (tylerd):

\[\frac{ (x+2)^2 }{ 25 } + \frac{ (y+5)^2 }{ 36 }= 1\]